OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)^3. for this do i add the two and 3 or do i multiply? also do i multiply the 8 by the 3?

OpenStudy (anonymous):

u multiply

OpenStudy (anonymous):

\[(8x^2)^3=8^3x^{2*3}=8^3x^6=512x^6\]

OpenStudy (campbell_st):

well you need split it 1st \[(8x^2)^3 = 8^3 \times (x^2)^3\] the power of a power rule for indices says \[(x^a)^b = x^{a \times b}\] this should help to simplify your question

OpenStudy (anonymous):

Choose the correct simplification of the expression d3 • d5. ok so here i add the exponents right?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz2)4. is this x^5y^5z^8?

OpenStudy (anonymous):

it is 4 where did u get 5 from?